Operations Report – Nov 6th

Crew 214 Operations Report 6-11-2019
SOL: 10
Name of person filing report: Andrew Wheeler

Non-nominal systems: Opportunity, Suit 10

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run
Hours run: 12h 40m
From what time last night: 1920
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 68%
Diesel Reading – 60%
Station Propane Reading – 80%
Ethanol Free Gasoline: N/A

Water (loft tank): 37 gallons
Water Meter: 1455889 units
Water (static tank): 223 gallons
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons

Toilet tank emptied: no

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 108.4
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 66%
Currently charging: yes

Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Curiosity rover used: yes
Hours: 114.3
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 74%
Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? N/A
CrewCar used and why, where? N/A

General notes and comments: The generator failed to recharge the batteries during the night.

Summary of internet: Functional
Summary of suits and radios: Nominal
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. 6 bags of soil obtained, which will be delivered to the Greenhab tomorrow.
Summary of Science Dome operations: All systems nominal. Sample processing following EVA was carried out and heating demonstration of hydrated minerals carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site